OpenStudy (anonymous):

2^5/2^3*4^-1/4^-2. Evaluate. Leave all your answers as powers using positive exponents.

OpenStudy (campbell_st):

well the rule for division is subtract the powers \[\frac{x^a}{x^b} = x^{a -b}\] and the rule for multiplication is add the powers \[x^a \times x^b = x^{a + b}\] and lastly power of a power... multiply the powers \[(x^a)^b = x^{a \times b}\] you need all 3 for this question and you need to write 4 as 2^2 so you have \[\frac{2^5}{2^3} \times \frac{(2^2)^{-1}}{(2^2)^{-2}} = \frac{2^5 \times 2^{-2}}{2^3 \times 2^{-4}}\] just just the rules above to simplify